Suppose that you generated a forecast with the some data, and you have the following errors:

Week (Et)
1
2 15.1
3 -2.8
4 1.2
5 -3.2
6 3.8
7

What is the MSE?

(Write your answer using ONE decimal place)
Answer #1

The following errons : Week (E+) (x) (y) 22 4 o 0 1 4 2 15.1 3012 -2.8 -8.4 1.2 16 4.8 3 4 5 6 7 -3.2. 25 -16 3.8 36 22.8 o 4Now we find the new y values: 5.3-0.8(1): 4.5 5.3 - 08(2) = 3.7 5.3-0.8(3) = 2.9 5.3 -0.8(4) = 2:1 5.3-0.8(5) - 1.3 5.3 -0.8